# Sub-second A-scan Acquisition Using Marginal Spectral-Domain Quantum Optical Coherence Tomography

### v1

@@ Abstract
We report an optimized implementation of spectral-domain quantum optical coherence tomography (SD-QOCT) capable of acquiring axial scans (A-scans) of multilayer samples without in the absence of mechanical scanning, at an unprecedented speed. Building on our group’s earlier work, we demonstrate a proof-of-concept system that integrates a diffraction grating, a high-resolution intensified CCD camera, and a high-flux photon-pair source operating in the VIS–NIR region ($\approx$800 nm). This configuration enables the acquisition of an entire marginal SD-QOCT interferogram in a single camera exposure, yielding a transverse A-scan with a record acquisition time of 100 ms and a penetration depth of $\approx$4 mm. The measured interferometric response shows excellent agreement with the theoretical model. These results represent a decisive step toward the practical deployment of SD-QOCT as a competitive imaging modality for biomedical applications.
@@ I Introduction
Quantum optical coherence tomography (QOCT) is the quantum analogue of optical coherence tomography (OCT) [Huang1991]. In QOCT, the classical low-coherence source is replaced by a source of entangled photon pairs; interference at a beamsplitter gives rise to the Hong–Ou–Mandel (HOM) effect [Hong1987]. Each interface within the sample generates a distinct HOM dip, enabling depth-resolved tomographic imaging [Abouraddy2002]. The use of frequency-correlated photon pairs endows QOCT with several advantages over classical OCT. For instance, when the SPDC pump bandwidth is sufficiently narrow, the technique becomes inherently immune to even-order dispersion in the sample, including group-velocity dispersion (GVD) [Steinberg1992, Franson1992]. Moreover, QOCT exhibits a quantum-enabled twofold enhancement in axial resolution compared with a classical OCT system of identical bandwidth [Nasr2003]. In fact, axial resolutions in the sub-micrometer regime have been demonstrated through careful engineering of the biphoton source [Okano2016, Katamadze2025].
@@
Several works have introduced creative strategies to overcome these limitations. For instance, full-field TD-QOCT [Ibarra2020] removes the need for transverse raster scanning on the $(x,y)$ plane, substantially reducing the time required to probe a three-dimensional object. Even more significantly, by spectrally resolving the TD-QOCT interferogram of a multilayer sample—transitioning from a purely time-domain function $P(\tau)$ to a hybrid frequency-resolved function $P(\tau,\omega_{s},\omega_{i})$—one can extract an A-scan without mechanical scanning. This approach, known as spectral-domain QOCT (SD-QOCT), has been demonstrated in the NIR spectral region (1550 nm), showing impressive performance and reducing A-scan acquisition times by orders of magnitude compared with TD-QOCT [Yepiz-Graciano2020]. Furthermore, as will be shown here, when the SPDC source is pumped with a sufficiently narrowband laser, the SD-QOCT scheme can be simplified, enabling even faster operation and single-point A-scan acquisition on sub-second timescales.
@@ II Theory
A standard TD-QOCT scheme directs the signal and idler photons from an SPDC source to the input ports of a beamsplitter (BS) after one photon has interacted with a multilayer sample and the other has been temporally delayed, producing the characteristic sequence of Hong–Ou–Mandel (HOM) dips [Hong1987] when monitoring the rate of coincidence events across the BS outputs. For an $N$-layer sample, the interferogram contains one HOM dip per physical interface, along with an additional set of $\binom{N}{2}$ artifact structures (dips or peaks) due to cross-interface interference that generally hinders accurate interpretation of the source morphology from the QOCT interferogram.
Whether an artifact associated with cross interference between two interfaces appears as a dip or a peak depends on the phase difference between the contributions from the two interfaces in question. Recent works have proposed effective strategies to mitigate these undesired artifact structures in QOCT implementations [Yepiz2023, Yepiz2025].
@@
By spectrally resolving the two photons in SD-QOCT, we retain the integrand of this last expression to yield the delay- and frequency-dependent interferogram $p_{c}(\tau,\omega_{s},\omega_{i})$, given by [Yepiz-Graciano2020]
@@
From this expression, we can integrate over $\omega_{s}$ to obtain the two-dimensional interferogram $p_{c}(\tau,\omega_{i})$, which can then be processed to extract information about the sample. In particular, there are two ways to obtain an A-scan from this interferogram: (i) tracing over the remaining frequency variable $\omega_{i}$ recovers the standard TD-QOCT interferogram $p_{c}(\tau)$ (the HOM dip), and (ii) fixing the temporal delay at $\tau=\tau_{0}$ yields the marginal interferogram $p_{c}(\tau_{0},\omega_{i})$ (see Fig. 1b). Applying a fast Fourier transform (FFT) to this marginal distribution produces the marginal temporal interferogram $p_{c}(\tau_{0},t)$, which directly reveals the A-scan information (see Fig. 1c) [Yepiz2019, KolenderskaFourierdomain2020].
@@ III Experiment
Reflection from PM leads to the SD-QOCT measurement stage, where the marginal interferogram $P_{c}(\tau_{0},\omega_{i})$ is obtained without the need for mechanical scanning. To achieve this, the idler photon is imaged through a $1.5\times$ telescope, formed by lenses $L_{8}$ ($f=500$ mm) and $L_{9}$ ($f=750$ mm), into the input port of an image-preserving optical delay line (OD) of approximately 28 m in length, providing a total propagation time of about 90 ns—sufficient to compensate for the internal delay of the ICCD camera (Andor iStar 334T) [Ibarra2020]. The image plane of the idler photon exiting the OD is subsequently reduced using a $5\times$ demagnifying telescope, which forms a focused image onto the vertical entrance slit of a monochromator (Andor Shamrock 500i), where a 1200-lines/mm, 750-nm-blaze diffraction grating disperses the spectral components. At the monochromator’s output plane, and in order to enhance the overall spectral resolution, a $5\times$ magnifying telescope—constructed with lenses $L_{10}$ ($f=50$ mm) and $L_{11}$ ($f=250$ mm)—projects the image onto the ICCD sensor. The ICCD is triggered by the detection of the signal photon in APDs; the transistor–transistor logic (TTL) pulse generated for each detection event is discriminated and appropriately delayed using a series of NIM electronic modules before reaching the camera.
@@ IV Results and discussion
From the measured interferogram and its corresponding FFT for the mirror sample, we can extract key performance parameters of our system, including the axial resolution and the effective imaging range. The axial resolution is given directly by the full width at half maximum (FWHM) of the HOM dip (Fig. 3e), which in our setup is approximately $500~\mu\text{m}$. This relatively modest resolution—compared with previous demonstrations by our group and others [Ibarra2020, Yepiz-Graciano2020, Yepiz-GracianoQuantumOptical2022, Kolenderska2025, Takeuchi2022]—stems from the narrow $\sim 1$ nm bandwidth of the photon pairs generated via collinear type-II SPDC in the PPKTP crystal.
Although this configuration was selected for the present proof-of-concept demonstration of the marginal SD-QOCT method due to its simplicity and ease of alignment, our system can readily accommodate photon-pair sources with substantially larger bandwidths. These include type-0 PPKTP crystals ($\sim 11$ nm) or even type-I BBO crystals ($\sim 100$ nm)—the latter being the source employed in our FF-QOCT demonstration, which held the record for the highest axial resolution in QOCT until 2022 [Ibarra2020, Takeuchi2022]. Future implementations of the M-SD-QOCT scheme will incorporate such broadband sources to achieve significantly enhanced axial resolution.
@@
Figure 4: a) SD-QOCT interferogram for the 1 mm glass slide, with the corresponding time-domain interferogram (FFT) shown in panel b. The sideplot in panel a illustrates a representative fixed-delay marginal interferogram, acquired with a 10 s ICCD exposure at the temporal-delay value indicated by the dashed green line; its FFT is displayed in the sideplot of panel b. Panels c and d present the corresponding numerical simulations. Panel e compares the TD-QOCT interferogram obtained from direct coincidence measurements (red dots), by integrating over frequency the experimental interferogram in panel a (black dots), and the theoretical interferogram in panel c (blue curve). The colored labels in panels b, d, and e correspond to the distinct reflective layers and artifact features of the sample (see text for details).
@@
Finally, Fig. 5a and Fig. 5b show the recorded marginal interferograms for the mirror and the glass-slide samples, acquired at temporal-delay positions of $c\tau_{0}=2$ mm for the mirror and $c\tau_{0}=3.8$ mm for the glass slide. These delay values satisfy the fixed-delay single exposure measurement criterion $c\tau_{0}\geq cT_{s}+c\sigma_{\tau}$, where in our case $c\tau_{0}=3.8$ $m$, $cT_{s}=3.2$ mm, and $c\sigma_{\tau}=0.5$ mm, all within the measured penetration depth of approximately 4 mm. The FFTs of these marginal interferograms are shown in Fig. 5c and Fig. 5d, corresponding to the reconstructed A-scans. The reflective surface of the mirror, as well as the front and rear interfaces of the glass slide, are clearly resolved at the expected optical-path differences based on the sample thickness. The blue curves in each panel show the corresponding numerical simulations, which agree exceptionally well with the experimental data.
@@
Importantly, as already mentioned, our system allows for a straightforward and significant upgrade: the axial resolution could be enhanced almost seamlessly to $\sim 11~\mu\text{m}$ by replacing the SPDC source with a type-0 PPKTP crystal [Kolenderska2025]. Additionally, although implementing artifact-removal strategies is beyond the scope of the present work, several of the techniques reported in the literature are fully compatible with our approach. These include phase-compensation methods that suppress artifacts via destructive interference, as well as optimization strategies based on genetic algorithms to identify and eliminate unwanted features in the interferogram [Yepiz2019, Yepiz2025]. We plan to incorporate such techniques in future optimized implementations of the M-SD-QOCT scheme.

### v2

@@ Abstract
We report an optimized implementation of spectral-domain quantum optical coherence tomography (SD-QOCT) capable of acquiring axial scans (A-scans) of multilayer samples in the absence of mechanical scanning, at an unprecedented speed. Building on our group’s earlier work, we demonstrate a proof-of-concept system that integrates a diffraction grating, a high-resolution intensified CCD camera, and a high-flux photon-pair source operating in the VIS–NIR region ($\approx$800 nm). This configuration enables the acquisition of an entire marginal SD-QOCT interferogram in a single camera exposure, yielding a transverse A-scan with a record acquisition time of 100 ms and a penetration depth of $\approx$4 mm. The measured interferometric response shows excellent agreement with the theoretical model. These results represent a decisive step toward the practical deployment of SD-QOCT as a competitive imaging modality for biomedical applications.
@@ I Introduction
Quantum optical coherence tomography (QOCT) is the quantum analogue of optical coherence tomography (OCT) [6]. In QOCT, the classical low-coherence source is replaced by a source of entangled photon pairs; interference at a beamsplitter gives rise to the Hong–Ou–Mandel (HOM) effect [5]. Each interface within the sample generates a distinct HOM dip, enabling depth-resolved tomographic imaging [1]. The use of frequency-correlated photon pairs endows QOCT with several advantages over classical OCT. For instance, when the SPDC pump bandwidth is sufficiently narrow, the technique becomes inherently immune to even-order dispersion in the sample, including group-velocity dispersion (GVD) [15, 2]. Moreover, QOCT exhibits a quantum-enabled twofold enhancement in axial resolution compared with a classical OCT system of identical bandwidth [13]. In fact, axial resolutions in the sub-micrometer regime have been demonstrated through careful engineering of the biphoton source [14, 8].
@@
Several works have introduced creative strategies to overcome these limitations. For instance, full-field TD-QOCT [7] removes the need for transverse raster scanning on the $(x,y)$ plane, substantially reducing the time required to probe a three-dimensional object. Even more significantly, by spectrally resolving the TD-QOCT interferogram of a multilayer sample—transitioning from a purely time-domain function $P(\tau)$ to a hybrid frequency-resolved function $P(\tau,\omega_{s},\omega_{i})$—one can extract an A-scan without mechanical scanning. This approach, known as spectral-domain QOCT (SD-QOCT), has been demonstrated in the NIR spectral region (1550 nm), showing impressive performance and reducing A-scan acquisition times by orders of magnitude compared with TD-QOCT [17]. Furthermore, as will be shown here, when the SPDC source is pumped with a sufficiently narrowband laser, the SD-QOCT scheme can be simplified, enabling even faster operation and single-point A-scan acquisition on sub-second timescales.
@@ II Theory
A standard TD-QOCT scheme directs the signal and idler photons from an SPDC source to the input ports of a beamsplitter (BS) after one photon has interacted with a multilayer sample and the other has been temporally delayed, producing the characteristic sequence of Hong–Ou–Mandel (HOM) dips [5] when monitoring the rate of coincidence events across the BS outputs. For an $N$-layer sample, the interferogram contains one HOM dip per physical interface, along with an additional set of $\binom{N}{2}$ artifact structures (dips or peaks) due to cross-interface interference that generally hinders accurate interpretation of the source morphology from the QOCT interferogram.
Whether an artifact associated with cross interference between two interfaces appears as a dip or a peak depends on the phase difference between the contributions from the two interfaces in question. Recent works have proposed effective strategies to mitigate these undesired artifact structures in QOCT implementations [12, 11].
@@
By spectrally resolving the two photons in SD-QOCT, we retain the integrand of this last expression to yield the delay- and frequency-dependent interferogram $p_{c}(\tau,\omega_{s},\omega_{i})$, given by [17]
@@
From this expression, we can integrate over $\omega_{s}$ to obtain the two-dimensional interferogram $p_{c}(\tau,\omega_{i})$, which can then be processed to extract information about the sample. In particular, there are two ways to obtain an A-scan from this interferogram: (i) tracing over the remaining frequency variable $\omega_{i}$ recovers the standard TD-QOCT interferogram $p_{c}(\tau)$ (the HOM dip), and (ii) fixing the temporal delay at $\tau=\tau_{0}$ yields the marginal interferogram $p_{c}(\tau_{0},\omega_{i})$ (see Fig. 1b). Applying a fast Fourier transform (FFT) to this marginal distribution produces the marginal temporal interferogram $p_{c}(\tau_{0},t)$, which directly reveals the A-scan information (see Fig. 1c) [3, 9].
@@ III Experiment
Reflection from PM leads to the SD-QOCT measurement stage, where the marginal interferogram $P_{c}(\tau_{0},\omega_{i})$ is obtained without the need for mechanical scanning. To achieve this, the idler photon is imaged through a $1.5\times$ telescope, formed by lenses $L_{8}$ ($f=500$ mm) and $L_{9}$ ($f=750$ mm), into the input port of an image-preserving optical delay line (OD) of approximately 28 m in length, providing a total propagation time of about 90 ns—sufficient to compensate for the internal delay of the ICCD camera (Andor iStar 334T) [7]. The image plane of the idler photon exiting the OD is subsequently reduced using a $5\times$ demagnifying telescope, which forms a focused image onto the vertical entrance slit of a monochromator (Andor Shamrock 500i), where a 1200-lines/mm, 750-nm-blaze diffraction grating disperses the spectral components. At the monochromator’s output plane, and in order to enhance the overall spectral resolution, a $5\times$ magnifying telescope—constructed with lenses $L_{10}$ ($f=50$ mm) and $L_{11}$ ($f=250$ mm)—projects the image onto the ICCD sensor. The ICCD is triggered by the detection of the signal photon in APDs; the transistor–transistor logic (TTL) pulse generated for each detection event is discriminated and appropriately delayed using a series of NIM electronic modules before reaching the camera.
@@ IV Results and discussion
From the measured interferogram and its corresponding FFT for the mirror sample, we can extract key performance parameters of our system, including the axial resolution and the effective imaging range. The axial resolution is given directly by the full width at half maximum (FWHM) of the HOM dip (Fig. 3e), which in our setup is approximately $500~\mu\text{m}$. This relatively modest resolution—compared with previous demonstrations by our group and others [7, 17, 16, 10, 4]—stems from the narrow $\sim 1$ nm bandwidth of the photon pairs generated via collinear type-II SPDC in the PPKTP crystal.
Although this configuration was selected for the present proof-of-concept demonstration of the marginal SD-QOCT method due to its simplicity and ease of alignment, our system can readily accommodate photon-pair sources with substantially larger bandwidths. These include type-0 PPKTP crystals ($\sim 11$ nm) or even type-I BBO crystals ($\sim 100$ nm)—the latter being the source employed in our FF-QOCT demonstration, which held the record for the highest axial resolution in QOCT until 2022 [7, 4]. Future implementations of the M-SD-QOCT scheme will incorporate such broadband sources to achieve significantly enhanced axial resolution.
@@
Figure 4: a) SD-QOCT interferogram for the 1 mm glass slide, with the corresponding time-domain interferogram (FFT) shown in panel b. The sideplot in panel a illustrates a representative fixed-delay marginal interferogram, acquired with a 10 s ICCD exposure at the temporal-delay value indicated by the dashed green line; its FFT is displayed in the sideplot of panel b. Panels c and d present the corresponding numerical simulations. Panel e compares the TD-QOCT interferogram obtained from direct coincidence measurements (red dots), by integrating over frequency the experimental interferogram in panel a (black dots), and the theoretical interferogram in panel c (blue curve). The colored labels in panels b, d, and e correspond to the distinct reflective layers and artifact features of the sample while the blue dashed line in all panels correspond to the temporal delay value used for capturing the measurements presented in Fig. 5 (see text for details).
@@
Finally, Fig. 5a and Fig. 5b show the recorded marginal interferograms for the mirror and the glass-slide samples, acquired at temporal-delay positions of $c\tau_{0}=1$ mm (blue dashed line in Fig. 3) for the mirror and $c\tau_{0}=3.8$ mm for the glass slide (blue dashed line in Fig. 4). These delay values satisfy the fixed-delay single exposure measurement criterion $c\tau_{0}\geq cT_{s}+c\sigma_{\tau}$, where in our case $c\tau_{0}=3.8$ $m$, $cT_{s}=3.2$ mm, and $c\sigma_{\tau}=0.5$ mm, all within the measured penetration depth of approximately 4 mm. The FFTs of these marginal interferograms are shown in Fig. 5c and Fig. 5d, corresponding to the reconstructed A-scans. The reflective surface of the mirror, as well as the front and rear interfaces of the glass slide, are clearly resolved at the expected optical-path differences based on the sample thickness. The blue curves in each panel show the corresponding numerical simulations, which agree exceptionally well with the experimental data.
@@
Importantly, as already mentioned, our system allows for a straightforward and significant upgrade: the axial resolution could be enhanced almost seamlessly to $\sim 11~\mu\text{m}$ by replacing the SPDC source with a type-0 PPKTP crystal [10]. Additionally, although implementing artifact-removal strategies is beyond the scope of the present work, several of the techniques reported in the literature are fully compatible with our approach. These include phase-compensation methods that suppress artifacts via destructive interference, as well as optimization strategies based on genetic algorithms to identify and eliminate unwanted features in the interferogram [3, 11]. We plan to incorporate such techniques in future optimized implementations of the M-SD-QOCT scheme.